OpenStudy (anonymous):

Hi guys, can you soulve this please? My Assignment... y-2x=3 3x+y=9 Substitution?

OpenStudy (saifoo.khan):

use eliminatiom

OpenStudy (saifoo.khan):

elimination

OpenStudy (anonymous):

--um...im just a beginner...? i need the coordinates?

OpenStudy (anonymous):

elimination y-y+-2x-3x=3-9 so -5x=-6 so x=6/5 then you can subts the x-value in either one of the equations to get y=27/5 hope i helped

OpenStudy (anonymous):

you can subtract the entire equations from each other, leaving 5x=6 (or x=6/5) and then put it in for x so y-12/5=3 which makes y "5and2/5"

OpenStudy (anonymous):

argh you beat me.

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

x = -3/11 y = 27/11

OpenStudy (anonymous):

--nice...i hope im like you.... thanks!

OpenStudy (anonymous):

khan please redo the problem

OpenStudy (anonymous):

saifoo, you're wrong. 27/11-9/11=18/11 which is not 9

OpenStudy (saifoo.khan):

i usd ny calc. :O

OpenStudy (saifoo.khan):

my*

OpenStudy (anonymous):

\[y=2x+3\] \[3x+2x+3=9\] \[5x=6\] \[x=\frac{6}{5}\]

OpenStudy (anonymous):

it lied. My calculator lies to me all the time. I have to send it to bed without dinner.

OpenStudy (anonymous):

lol jibble you crack me up

OpenStudy (anonymous):

\[3\times \frac{6}{5}+y=9\] \[frac{18}{5}+y=9\] \[y=9-\frac{18}{5}\] \[y=\frac{27}{5}=5.4\]

OpenStudy (anonymous):

you can just remove its battery and wait for it to come and apologise

OpenStudy (anonymous):

I'd probably be waiting for a while, huh?

OpenStudy (anonymous):

lol ya while you wait it will be learning its lesson